OpenStudy (anonymous):

How would I find the inverse function (if it exists) of f(x) =ln(1 - 6 x)?

OpenStudy (anonymous):

\[f(x)=\ln(x)\]Is the inverse function of:\[g(x)=e^x\]Therefore,\[f(x)=\ln(1-6x)\]Is the inverse function of:\[g(x)=e^{1-6x}\]

OpenStudy (anonymous):